OpenStudy (anonymous):

Using absolute value inequalities. Confused. Quiz Question: A furniture maker uses the specification 19.88 (less than or equal to) w (less than or equal to) 20.12 for the width w in inches of a desk drawer. Write the specification as an absolute value inequality. My (wrong) Answer: |w - 0.12| (less than or equal to) 20 To be honest, I guess on that question because I didn't know what I was supposed to be doing. The Right Answer Was: |w - 20| (less than or equal to) 0.12 My teacher then told me to look at a problem in our book that only confused me further

OpenStudy (anonymous):

In car racing, a car must meet specific dimensions to enter a race. Officials use a template to ensure these specifications are met. What absolute value inequality describes heights of the model of race car shown within the indicated tolerance. 53-51/2 = 2/2 = 1 Find the tolerance. -1 (less than or equal to) h - 52 (less than or equal to) 1 Use h for the height of the race car. Write a compound inequality. |h-52| (less than or equal to) 1